OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4

OpenStudy (anonymous):

Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

1)y-3=4(x+6) 2)y-3=4x+24 y=4x+27